# Mock Eisenstein series associated to partition ranks

### v1

@@ Abstract.
In this paper, we introduce a new class of mock Eisenstein series, describe their modular properties, and write the partition rank generating function in terms of so-called partition traces of these. Moreover, we show the Fourier coefficients of the mock Eisenstein series are integral and we obtain a holomorphic anomaly equation for their completions.
@@ 1. Introduction and statement of results
[…]
@@
Dyson conjectured that reducing the rank (mod $5$) (resp. $7$) divides the partitions of $45n+$ 4 (resp. $57n+$ 5) into $5$ (resp. $7$) sets of equal size. This conjecture was proven by Atkin and Swinnerton-Dyer [5]. In the same paper, Dyson also conjectured the existence of another statistic, which he called the “crank” and which should explain all three partition congruences. Garvan [12] found a crank for vector partitions and Andrews–Garvan [4] defined a crank for ordinary partitions. Letting $o(\lambda)$ denote the number of ones in a partition $\lambda$, and $\mu(\lambda)$ the number of parts strictly larger than $o(\lambda)$, the crank is defined as
@@
(see [5]) where $(a)_{n}=(a;q)_{n}:=\prod_{j=0}^{n-1}(1-aq^{j})$ for $n\in\mathbb{N}_{0}\cup\{\infty\}$. Let $M(m,n)$ denote the number of partitions of $n$ with crank $m$, except for [^1] The correct combinatorial values for the anomalous case $n=1$ are $1$ if $m=0$ and 00 for $m\neq 0$. $n=1$ where $M(-1,1)=-M(0,1)=M(1,1):=1$ as given by the following generating function [4]
$C(\zeta;q):=\sum_{\begin{subarray}{c}m\in\mathbb{Z}\\ n\geq 0\end{subarray}}M(m,n)\zeta^{m}q^{n}=\frac{(q)_{\infty}}{(\zeta q)_{\infty}\left(\zeta^{-1}q\right)_{\infty}}.$ (1.2)
Modularity properties of the rank and of the crank generating function differ significantly: The crank generating function is basically a meromorphic Jacobi form, whereas the rank generating function is a “mock Jacobi form” [8]: it only transforms like a Jacobi form after adding a non-holomorphic term (see Section 2.3 for the precise transformations of the rank generating function). We next consider the crank moments [3]
$C_{k}(q):=\sum_{n\geq 0}\sum_{m\in\mathbb{Z}}m^{k}M(m,n)q^{n}.$
These moments can be expressed in terms of quasimodular forms. In [1], the authors expressed these as a so-called partition Eisenstein trace. Here, for a sequence of functions $h=\{h_{k}\}_{k\in\mathbb{N}}$, define, for $n\in\mathbb{N}$, the $n$-th partition trace with respect to $h$ and a function $\phi$ on partitions as
@@
$G_{k}(\tau):=-\frac{B_{k}}{2k}+\sum_{m,n\geq 1}m^{k-1}q^{mn}\qquad(q:=e^{2\pi i\tau},\,\tau\in\mathbb{H}:=\{w\in\mathbb{C}:\operatorname{Im}(w)>0\}),$
@@
It is natural to wonder whether a result like Theorem 1.1 involving the rank moments
$R_{k}(q):=\sum_{n\geq 0}\sum_{m\in\mathbb{Z}}m^{k}N(m,n)q^{n}.$ (1.5)
exists. In this paper, we show that this is indeed the case. As mentioned above, the rank moments are related to mock modular forms. Hence, they do not admit an expansion as a trace of modular Eisenstein series like crank moments. As our first result, we write the rank moments as traces of mock Eisenstein series $f_{k}$ defined in Subsection 3.1. The aim of this paper is to understand these functions $f_{k}$. We call a real-analytic function $f^{*}(\tau,\overline{\tau})$ a quasi-completion of $f(\tau)$, if $f^{*}(\tau,\overline{\tau})$ transforms like a quasimodular form (see (2.2)) and if [^3] Here and throughout, we consider $\tau$ and $\overline{\tau}$ as independent variables. $\lim_{\overline{\tau}\to-i\infty}$ $f^{*}(\tau,\overline{\tau})=f(\tau)$. If a quasi-completion $f^{*}(\tau,\overline{\tau})$ transforms as a modular form, then we call $f^{*}(\tau,\overline{\tau})$ a completion. If it is clear from the context, then we also just write $f^{*}(\tau)$ instead of $f^{*}(\tau,\overline{\tau})$. Our first result is the following theorem.
@@
    The function $f_{k}$ has a quasi-completion [^4] We consider quasi-completions, because they are more analogous to the functions in Theorem 1.1. In the sequel, we also study the completions $\widehat{f_{k}}$ of $f_{k}$. $f^{*}_{k}$ which satisfies, for $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\mathrm{SL}_{2}(\mathbb{Z})$,
@@
Remark 1.3.
(1) As for the Eisenstein series, we have $f_{k}=0$ if $k$ is odd; see Remark 3.1.
(2) The transformation of $f_{k}^{*}$ agrees with that of $G_{k}$ up to a factor of $3$ in front of $c\tau+d$. This is explained by the (mock) Jacobi forms underlying the rank and crank statistics: namely, for the rank the corresponding index is $-\frac{3}{2}$ and for the crank it is $-\frac{1}{2}$.
@@
$g_{\ell}(\tau):=\left(1-2^{\ell-1}\right)\frac{B_{\ell}}{2\ell}+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\end{subarray}}(2n-3m)^{\ell-1}q^{mn}-\sum_{\begin{subarray}{c}n-1\geq 6m\geq 6\end{subarray}}(n-6m)^{\ell-1}q^{mn}.$ (1.7)
For $r\in\mathbb{N}$, the $r$-th Fourier coefficient of $g_{\ell}$ is a polynomial in some of the divisors of $r$: namely those positive divisors $m,n$ satisfying the inequality $2n-1\geq 3m$ or $n-1\geq 6m$. Hence, our functions are in spirit of the “mock Eisenstein series” studied by Zagier [18, p.15] and Mertens–Ono–Rolen [13, equation (1.4)].
@@
Finally, in part (1) of the next theorem, we give an explicit formula for $D(f_{k})$ and explain how the raising and lowering operators act on the completions $\widehat{f}_{k}$ of the [^5] See Subsection 2.1 for the definition; in particular, $\widehat{f}_{k}$ and $f_{k}^{*}$ agree if $k\neq 2$. $f_{k}$. More precisely, let ^$\widehat{\mathcal{F}}:=\mathbb{C}[\widehat{f}_{2},\widehat{f}_{4},\widehat{f}_{6},\ldots,\widehat{G}_{2},G_{4},G_{6}]$ and define the raising and the lowering operator by
@@
In parts (2) and (3), we obtain a recursive expression for the action of the raising and lowering operators on $\widehat{f}_{k}$. In the physics literature, the latter of these falls into the realm of holomorphic anomaly equations. We let $\delta_{\mathcal{S}}:=1$ if a statement $\mathcal{S}$ holds and $\delta_{\mathcal{S}}:=0$ otherwise.
@@
    If $\widehat{f}\in\widehat{\mathcal{F}}$ is of weight $k$, then $(D+\frac{2k}{3}f_{2}^{*})(\widehat{f})$ transforms modular of weight $k+2$.
Remark 1.7.
We expect that the algebra $\mathcal{F}$ is freely generated by the $f_{k}$ for $k$ even and $G_{2},G_{4}$, and $G_{6}$ (see also Question (2) in Section 6). In particular, this would imply that the subalgebra $\mathbb{Q}[f_{2},f_{4},\ldots]$ of $\mathcal{F}$ is not closed under the derivation $D$, since Theorem 1.6 (1) involves the modular Eisenstein series.
@@ Acknowledgements
We thank Caner Nazaroglu and Kilian Rausch for helpful comments on a previous version of the manuscript. The first and the second authors were funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001179) and the first and the third authors by the SFB/TRR 191 “Symplectic Structure in Geometry, Algebra and Dynamics”, funded by the DFG (Projektnummer 281071066 TRR 191).
@@ 2.1. Modular forms and quasimodular forms
    We have [^6] The $f_{j}$ here should not be confused with the “mock Eisenstein series” $f_{j}$.
    $f(\tau)=\sum_{j=0}^{s}\frac{f_{j}(\tau)}{v^{j}},$
    for some holomorphic functions $f_{j}:\mathbb{H}\to\mathbb{C}$ with $f_{s}\neq 0$.
@@
By convention, the zero function is an almost holomorphic modular form of depth $\infty$- ∞.
@@
More concretely, for a quasimodular form $g$ of weight $k$ and depth $s$ there exist holomorphic functions $g_{j}:\mathbb{H}\to\mathbb{C}$ for $j\in\{0,\ldots,s\}$ with $g_{0}=g$ such that
$(c\tau+d)^{-k}g\!\left(\frac{a\tau+b}{c\tau+d}\right)\>=\>\sum_{j=0}^{s}g_{j}(\tau)\Bigl{(}\frac{c}{c\tau+d}\Bigr{)}^{j}$
for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$. Similarly, we say an analytic function $g^{*}(\tau,\overline{\tau})$ transforms like a quasimodular form if there exist real-analytic functions $g_{j}^{*}(\tau,\overline{\tau})$ so that
$(c\tau+d)^{-k}g^{*}\!\left(\frac{a\tau+b}{c\tau+d},\frac{a\overline{\tau}+b}{c\overline{\tau}+d}\right)\>=\>\sum_{j=0}^{s}g_{j}(\tau,\overline{\tau})\Bigl{(}\frac{c}{c\tau+d}\Bigr{)}^{j}$ (2.2)
for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$. The space of all quasimodular forms is a free algebra with generators $G_{2},G_{4}$, and $G_{6}$. It follows from Ramanujan’s differential equations, [14, equations (1), (2)]
@@
that this algebra is closed under differentiation. We also require the Serre derivative (see, e.g., [17, p. 48])
@@ 2.2. Crank moments
Set $\zeta:=e^{2\pi iz}$. By (1.2) and [16, equation (7)], the crank generating function can be expressed in terms of the Eisenstein series. Recall that $G_{k}=0$ if $k$ is odd.
@@ 2.3. Mock modularity of the rank generating function
$R^{\#}(z;\tau):=\left(\frac{R(\zeta;q)}{\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}}q^{-\frac{1}{24}}+\frac{1}{2}q^{-\frac{1}{6}}\sum_{\pm}\pm\zeta^{\mp 1}S(3z\pm\tau;3\tau)\right)e^{12\pi^{2}G_{2}(\tau)z^{2}},$ (2.9)
@@
[…]

### v2

# Eisenstein–type series associated to partition ranks

@@ Abstract.
In this paper, we introduce a class of functions that behave like classical Eisenstein series in many ways, but with a key distinction: only their non-holomorphic completions transform like (quasi)modular forms. We show how the partition rank generating function can be expressed in terms of partition traces of these functions. A key feature of our construction is that the completions satisfy a holomorphic anomaly equation–a phenomenon typically seen in the context of quantum field theory and string theory. We also show that the Fourier coefficients of these Eisenstein-type series are integral.
@@ 1. Introduction and statement of results
In this paper, we introduce a new class of Eisenstein-type series. These series share several structural features with classical Eisenstein series, but differ in a crucial respect: only their non-holomorphic completions transform as (quasi)modular forms. To motivate their definition, we first explore their connection to the rank generating function, before examining their further properties in detail.
1.1. Background
@@
$p(5n+4)\equiv 0\ \,(\mathrm{mod}\,5),\quad p(7n+5)\equiv 0\ \,(\mathrm{mod}\,7),\quad p(11n+6)\equiv 0\ \,(\mathrm{mod}\,11).$
@@
Dyson conjectured that reducing the rank (mod $5$) (resp. $7$) divides the partitions of $5n+4$ (resp. $7n+5$) into $5$ (resp. $7$) sets of equal size. This conjecture was proven by Atkin and Swinnerton-Dyer [5]. In the same paper, Dyson also conjectured the existence of another statistic, which he called the “crank” and which should explain all three partition congruences. Garvan [15] found a crank for vector partitions and Andrews–Garvan [4] defined a crank for ordinary partitions. Letting $o(\lambda)$ denote the number of ones in a partition $\lambda$, and $\mu(\lambda)$ the number of parts strictly larger than $o(\lambda)$, the crank is defined as
@@
(see [5]) where $(a)_{n}=(a;q)_{n}:=\prod_{j=0}^{n-1}(1-aq^{j})$ for $n\in\mathbb{N}_{0}\cup\{\infty\}$. Let $M(m,n)$ denote the number of partitions of $n$ with crank $m$, except for [^1] The correct combinatorial values for the anomalous case $n=1$ are $1$ if $m=0$ and 0 for $m\neq 0$. $n=1$ where M(1,1)=M(0,1)=M(1,1):=1M(-1,1)=-M(0,1)=M(1,1):=1 as given by the following generating function [4]
$C(\zeta;q):=\sum_{\begin{subarray}{c}n\geq 0\\ m\in\mathbb{Z}\end{subarray}}M(m,n)\zeta^{m}q^{n}=\frac{(q)_{\infty}}{(\zeta q)_{\infty}\left(\zeta^{-1}q\right)_{\infty}}.$ (1.2)
Modularity properties of the rank and of the crank generating function differ significantly: the crank generating function is basically a meromorphic Jacobi form, whereas the rank generating function is a “mock Jacobi form” [8]: it only transforms like a Jacobi form after adding a non-holomorphic term (see Subsection 2.3 for the precise transformations of the rank generating function). We next consider the crank moments [3]
$C_{k}(q):=\sum_{\begin{subarray}{c}n\geq 0\\ m\in\mathbb{Z}\end{subarray}}m^{k}M(m,n)q^{n}.$
These moments can be written in terms of quasimodular forms. In [1], the authors expressed these as a so-called partition Eisenstein trace. Here, for a sequence of functions $h=\{h_{k}\}_{k\in\mathbb{N}}$, define, for $n\in\mathbb{N}$, the $n$-th partition trace with respect to $h$ and a function $\phi$ on partitions as
@@
$G_{k}(\tau):=-\frac{B_{k}}{2k}+\sum_{n,m\geq 1}m^{k-1}q^{nm}\qquad(q:=e^{2\pi i\tau},\,\tau\in\mathbb{H}:=\{w\in\mathbb{C}:\operatorname{Im}(w)>0\}),$
@@
1.2. Main results
A result like Theorem 1.1 involving the rank moments
$R_{k}(q):=\sum_{\begin{subarray}{c}n\geq 0\\ m\in\mathbb{Z}\end{subarray}}m^{k}N(m,n)q^{n}$ (1.5)
may give rise to an interesting generalization of the Eisenstein series. In this paper, we show that this is indeed the case. We obtain a surprisingly simple definition of Eisenstein-type series [^3] We emphasize that these functions are not constructed by averaging, as is done for classical Eisenstein series. They are called Eisenstein-type because they satisfy properties analogous to those of classical Eisenstein series. $g_{\ell}$ in (1.7) which seem not to be studied in the literature before. As mentioned above, the rank moments are related to mock modular forms. Hence, they do not admit an expansion as a trace of modular Eisenstein series like crank moments. As our first result, we write the rank moments as traces of the Eisenstein–type series $f_{k}$ defined in Subsection 3.1. The aim of this paper is to understand these functions $f_{k}$. We call a real-analytic function $f^{*}(\tau,\overline{\tau})$ a quasi-completion of $f(\tau)$, if $f^{*}(\tau,\overline{\tau})$ transforms like a quasimodular form (see (2.2)) and if [^4] Here and throughout, we consider $\tau$ and $\overline{\tau}$ as independent variables. $\lim_{\overline{\tau}\to-i\infty}$ $f^{*}(\tau,\overline{\tau})=f(\tau)$. If a quasi-completion $f^{*}(\tau,\overline{\tau})$ transforms as a modular form, then we call $f^{*}(\tau,\overline{\tau})$ a completion. If it is clear from the context, then we also just write $f^{*}(\tau)$ instead of $f^{*}(\tau,\overline{\tau})$. Our first result is the following theorem.
@@ 1.2. Main results
    The function $f_{k}$ has a quasi-completion [^5] We consider quasi-completions, because they are more analogous to the functions in Theorem 1.1. In the sequel, we also study the completions $\widehat{f_{k}}$ of $f_{k}$; see (3.4) and (3.6) for their definitions. $f^{*}_{k}$ which satisfies, for $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\mathrm{SL}_{2}(\mathbb{Z})$,
@@
Remark.
  1. (1)
    As for the Eisenstein series, we have $f_{k}=0$ if $k$ is odd; see Remark 3.1.
  2. (2)
    The functions $f_{k}$ and $f_{k}^{*}$ are of intrinsic interest, as they are a step up in complexity compared to quasimodular forms.
  3. (3)
    The transformation of $f_{k}^{*}$ agrees with that of $G_{k}$ up to a factor of $3$ in front of $c\tau+d$. This is explained by the (mock) Jacobi forms underlying the rank and crank statistics: namely, for the rank the corresponding index is $\frac{3}{2}$ and for the crank it is $\frac{1}{2}$.
@@
$g_{\ell}(\tau):=\left(1-2^{\ell-1}\right)\frac{B_{\ell}}{2\ell}+\sum_{\begin{subarray}{c}2n-1\geq 3m\geq 3\end{subarray}}(2n-3m)^{\ell-1}q^{nm}-\sum_{\begin{subarray}{c}n-1\geq 6m\geq 6\end{subarray}}(n-6m)^{\ell-1}q^{nm}.$ (1.7)
For $r\in\mathbb{N}$, the $r$-th Fourier coefficient of $g_{\ell}$ is a polynomial in some of the divisors of $r$: namely those positive divisors $n,m$ satisfying the inequality $2n-1\geq 3m$ or $n-1\geq 6m$. Hence, the $g_{k}$ are in spirit of the “mock Eisenstein series” studied by Zagier [25, p.15] and Mertens–Ono–Rolen [19, equation (1.4)]. Moreover, compare $g_{\ell}$ with
$F_{\ell}^{[2]}(\tau):=\sum_{\begin{subarray}{c}n>m\geq 1\\ n-m\text{ odd }\end{subarray}}(-1)^{n}m^{\ell-1}q^{\frac{nm}{2}}$
which can be expressed as the Rankin–Cohen bracket of a mock modular form and $\eta^{3}$ as explained in [10, Example 4 on p. 44]. This function goes back to [12, 13], where the authors studied characters of the $\mathcal{N}=4$ superconformal algebra in two dimensions and the elliptic genus of K3 surfaces.
@@
As a direct corollary, we see that the algebra $\mathcal{G}:=\mathbb{Q}[g_{2},g_{4},\ldots,G_{2},G_{4},\ldots]$ is closed under $D$, and that the elements of $\mathcal{G}$ admit a quasi-completion:
Corollary 1.4.
We have $\mathcal{G}=\mathcal{F}$, where the algebra $\mathcal{F}$ is given by Theorem 1.2 (3).
@@
Finally, in part (1) of the next theorem, we give an explicit formula for $D(f_{k})$ and explain how the raising and lowering operators act on the completions $\widehat{f}_{k}$ of the [^6] See Subsection 2.1 for the definition; in particular, $\widehat{f}_{k}$ and $f_{k}^{*}$ agree if $k\neq 2$. $f_{k}$. More precisely, let $\widehat{\mathcal{F}}:=\mathbb{C}[\widehat{f}_{2},\widehat{f}_{4},\widehat{f}_{6},\ldots,\widehat{G}_{2},G_{4},G_{6}]$ and define the raising and the lowering operator by
@@
In parts (2) and (3), we obtain a recursive expression for the action of the raising and lowering operators on $\widehat{f}_{k}$. In physics literature, the latter of these falls into the realm of “holomorphic anomaly equations”. In string theory, mirror symmetry, enumerative geometry, and topological field theory, a holomorphic anomaly refers to the phenomenon where a generating function (or its coefficients), instead of being purely holomorphic, acquires a controlled dependence on both $\tau$ and $\overline{\tau}$. A holomorphic anomaly equation describes this controlled dependency through a recursive relation that expresses the anti-holomorphic variation in terms of functions of lower genus or weight. While these may themselves carry non-holomorphic dependence, the anomaly equation organizes this structure in a precise and constrained way (see for example [16, 17, 18, 20]). We let $\delta_{\mathcal{S}}:=1$ if a statement $\mathcal{S}$ holds and $\delta_{\mathcal{S}}:=0$ otherwise.
@@
[…]
Remark.
We expect that the algebra $\mathcal{F}$ is freely generated by the $f_{k}$ for $k$ even and $G_{2},G_{4}$, and $G_{6}$ (see also question (2) in Section 6). In particular, this would imply that the subalgebra $\mathbb{Q}[f_{2},f_{4},\ldots]$ of $\mathcal{F}$ is not closed under applying $D$, since Theorem 1.6 (1) involves the modular Eisenstein series.
1.3. Structure of the paper
@@ Acknowledgements
We thank Caner Nazaroglu, Kilian Rausch, and Don Zagier for helpful comments on a previous version of the manuscript. The first and the second authors were funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101001179) and the first and the third authors by the SFB/TRR 191 “Symplectic Structure in Geometry, Algebra and Dynamics”, funded by the DFG (Projektnummer 281071066 TRR 191).
@@ 2.1. Modular forms and quasimodular forms
    We have, [^7] The $f_{j}$ here should not be confused with the Eisenstein–type series $f_{j}$. for some holomorphic functions $f_{j}:\mathbb{H}\to\mathbb{C}$ with $f_{s}\neq 0$,
    $f(\tau)=\sum_{j=0}^{s}\frac{f_{j}(\tau)}{v^{j}}.$
@@
By convention, the function zero is an almost holomorphic modular form of depth -$\infty$.
@@
More concretely, for a quasimodular form $g$ of weight $k$ and depth $s$ there exist holomorphic functions $g_{j}:\mathbb{H}\to\mathbb{C}$ for $j\in\{0,\ldots,s\}$ with $g_{0}=g$ such that, for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$,
$(c\tau+d)^{-k}g\!\left(\frac{a\tau+b}{c\tau+d}\right)\>=\>\sum_{j=0}^{s}g_{j}(\tau)\Bigl(\frac{c}{c\tau+d}\Bigr)^{j}.$
Similarly, we say an analytic function $g^{*}(\tau,\overline{\tau})$ transforms like a quasimodular form if there exist real-analytic functions $g_{j}^{*}(\tau,\overline{\tau})$ so that, for all $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{Z})$,
$(c\tau+d)^{-k}g^{*}\!\left(\frac{a\tau+b}{c\tau+d},\frac{a\overline{\tau}+b}{c\overline{\tau}+d}\right)\>=\>\sum_{j=0}^{s}g_{j}(\tau,\overline{\tau})\Bigl(\frac{c}{c\tau+d}\Bigr)^{j}.$ (2.2)
The space of all quasimodular forms is a free algebra with generators $G_{2},G_{4}$, and $G_{6}$. It follows from Ramanujan’s differential equations, [21, equations (1), (2)]
@@
that this algebra is closed under differentiation. We also require the Serre derivative (see, e.g., [24, p. 48])
@@ 2.2. Crank moments
Set $\zeta:=e^{2\pi iz}$. By (1.2) and [23, equation (7)], the crank generating function can be expressed in terms of the Eisenstein series. Recall that $G_{k}=0$ if $k$ is odd.
@@ 2.3. Mock modularity of the rank generating function
$\hskip-5.69046ptR^{\#}(z;\tau):=\left(\frac{R(\zeta;q)}{\zeta^{\frac{1}{2}}-\zeta^{-\frac{1}{2}}}q^{-\frac{1}{24}}+\frac{1}{2}q^{-\frac{1}{6}}\sum_{\pm}\pm\zeta^{\mp 1}S(3z\pm\tau;3\tau)\right)e^{12\pi^{2}G_{2}(\tau)z^{2}},$ (2.9)
@@
with $E(y):=2\int_{0}^{y}e^{-\pi t^{2}}dt$. She showed, building on work of Zwegers [26], the following transformation: